I was wondering if you could help me with this problem. I need it as soon as possible. Thank you! A train averaged 120 km/h for the first two thirds of a trip and 100 km/h for the whole trip. Find its average speed for the last third of the trip. I think the answer is 75km/h but i don't know how to set it up. Thank you.
Answer by venugopalramana(3286)

A train averaged 120 km/h for the first two thirds of a trip and 100 km/h for the whole trip. Find its average speed for the last third of the trip. I think the answer is 75km/h but i don't know how to set it up. Thank you.
LET THE TRIP BE FOR SAY X KM.
2/3 RD. OF TRIP =X*2/3=2X/3 KM.....IT TRAVELLED AT 120 KMPH
SO TIME TAKEN = 2X/(3*120)=X/180 HRS.
WHOLE TRIP AVERAGE SPEED =100 KMPH ..SO TIME TAKEN FOR WHOLE TRIP = X/100 HRS.
SO TIME TAKEN FOR 1/3 RD. OF TRIP=X/100 - X/180 =X(180-100)/18000=X/225 HRS.
DISTANCE TRAVELLED =X*1/3=X/3
SO AVERAGE SPEED FOR THIS PART = (X/3)/(X/225)=225/3 =75 KMPH